# Studying the GRAiNITA concept: first test beam results

### v1

@@ 1 Introduction
A small-scale GRAiNITA prototype, employing millimetric grains of ZnWO4 crystals, was constructed to perform an initial validation of the detector concept. Experimental results confirmed that the scintillation signal remains localized near its production point and that a substantial number of photo-electrons are generated [Barsuk:2023uva], thereby providing strong evidence supporting the viability of the GRAiNITA design. In particular, measurements using several hundred cosmic muons yielded approximately 400 photo-electrons for an estimated deposited energy of about 40 MeV, indicating the potential to achieve an energy resolution contribution from photo-electron statistics at the level of . Simulation studies using a simplified set-up with GEANT4 [G4] have shown that the 1% term worsens to about 2% when including the sampling fluctuations [bib:IEEE2021]. This term is smaller than in usual sampling calorimeter due to the smallness of the grains and the high density of the scintillator.
@@ 2 Experimental setup
The 16-channels GRAiNITA small-size prototype is shown in figure 1. It consists of a metallic container with an active volume of mm3 and it is equipped with 16 O-2(200) WLS fibers (1 mm diameter ) arranged in square geometry and inter-spaced by 7 mm. In order to monitor the behaviour of the setup, a clear fiber, unpolished along 1 cm, is placed in the middle of the WLS ones and connected to a green LED222Broadcom HLMP-CM1A-560DD, wavelength from 500 to 540 nm. to inject the light in the active volume.
Data-taking took place during two days at the SPS beam (H2 line) delivering muons and pion secondary beams. In total about 2 million of muons events and 48 millions of pions events were recorded. Given the smallness of the GRAiNITA prototype, the estimated interaction probability for the pions is approximately 20%. The data acquisition was divided into two periods. In the first period (P1), ZnWO4 grains were immersed in water, while in the second period (P2) they were immersed in a water-based sodium polytungstate solution, hereafter referred to as the heavy liquid, characterized by a refractive index of 1.5 and a density of . For both periods, data were collected using pion and muon beams. Toward the end of the acquisition, partial leaks of the heavy liquid were observed due to a container closure not completely liquid-tight. Consequently, we chose to focus on the data obtained with water to evaluate non-uniformity effects, while using the heavy-liquid data primarily for calibration purposes, since no significant differences in non-uniformity are expected between the two liquid configurations. The experimental setup used is shown in figure 2 and consists of a set of 3 wire chambers (DWC) for reconstructing the beam tracks, 2 sets of dual scintillator for the triggering surrounding the GRAiNITA prototype.
Each of the 16 WLS fibers of GRAiNITA is coupled to a SiPM mounted on a PCB for supply, gain stability and amplification. The signals from the SiPMs are digitized at 3.2 GigaSample/s with a WaveCatcher module [bib:WaveC]. The number of photo-electrons is recorded in a 25 s windows corresponding to the scintillating decay time constant of the grains.
@@ 3.3 Light confinement
Using the largest statistics run with pions from P2 and the four more central fibres, the average number of photo-electrons read by a given fibre as a function of the position of the track with respect to this fibre is shown in figure 4. These observations clearly confirm that the scintillation light remains confined near its point of production. The right-hand plot displays the dependence of the mean number of photo-electrons on the distance from the fibre centre. Approximately 50% of the detected light originates from particle interactions occurring within 2.7 mm from the fibre centre, corresponding to slightly less than half the spacing between adjacent fibres.
The results described in sections 3.2 and 3.3 are used to assess more precisely the fibre position. It is determined via the minimization of the sum of the number of photo-electrons in a square of bins in a bin windows around the expected position. The minimum in both directions is taken as the actual fibre position. The precision of this method is about the size of a bin (m).
3.4 Assessment of the stochastic term of the energy resolution
An encouraging number of approximately 400 photo-electrons was obtained with the cosmic muons [Barsuk:2023uva] leading to a stochastic term on the energy resolution of the order of 1%. The much larger data sample recorded during the test beam and the possibility to better define the tracks used for the measurement calls for a new estimate. In order to avoid border effects, only the events where tracks are hitting the central region of the prototype are used, and the data recorded with muon beams is analysed. A Landau distribution convoluted with a Gaussian function as described in section 4.2 is fitted to the distribution of the total number of photo-electrons recorded on the 16 WLS fibres. The fit for the two periods is shown in figure 5.
The number of photoelectrons measured for water and for the heavy liquid are 479 and 452, respectively, and are in very good agreement with the values obtained using cosmic muons. [Barsuk:2023uva].
@@ 5.3 Impact of the non-uniformity response on the constant term
A simplified model of a full scale calorimeter module is simulated using Geant4 Monte-Carlo [G4]. The size of the calorimeter module is mm3. The depth of 400 mm corresponds to 25 radiation lengths (), which is sufficient to fully contain electromagnetic showers initiated by 25 GeV photons. This volume contains a grid of 24 by 24 WLS fibres evenly distributed in the corners of squares of 7 by 7 mm2, to reproduce the small-scale GRAiNITA prototype. The WLS fibres are simulated as cylinders of 1 mm diameter. In order to assess in a reasonable amount of CPU time the impact of the measured non-uniformity the grains are not simulated individually but a single material was simulated with partial contents of the scintillator and heavy liquid corresponding to the ones measured with the physical detector used for the beam test. The mass fractions of the detector material simulated are : 20.77 % of heavy liquid and 79.23 % of ZnWO4. The radiation length, calculated with the Geant4 framework, of this mixture of materials is 1.64 cm which aligns with manual cross-calculations.
@@ 6 Conclusion
Data from a two-day period in June 2024 using the H9 test beam of the SPS at CERN recorded with the small-scale GRAiNITA prototype is analyzed. These measurements confirm that the contribution to the energy resolution due to the number of photo-electrons is of the order of . With preliminary simulations of the grains and the heavy liquid [bib:IEEE2021], this term goes to about . More importantly, the data recorded during this test beam period enables a first estimation of the constant term associated with detector non-uniformity. Despite the limitations of these measurements due to the small size of the detector and the use of pions, the constant term due to the non-uniformity should be significantly below 1%.

### v2

@@ 1 Introduction
A small-scale GRAiNITA prototype, employing millimetric grains of ZnWO4 crystals, was constructed to perform an initial validation of the detector concept. Indeed, ZnWO4 is a well suited non-hygroscopic inorganic scintillator material for GRAiNITA of high density 7.87 g/cm3), effective atomic number (61) with a good light yield ( 9000 photoelectrons /MeV) [4, 6] and grains can be obtained via spontaneous crystallization from a flux melt [9]. Measurements using several hundred cosmic muons yielded approximately 400 photo-electrons for an estimated deposited energy of about 40 MeV [2], indicating the potential to achieve an energy resolution contribution from photo-electron statistics at the level of . Simulation studies with a simplified GEANT4 set-up [1] indicate that the 1% term degrades to approximately 2% [7] as a result of event-to-event sampling fluctuations in the partition of deposited energy between the scintillator grains, which contribute to the detected signal, and the inter-grain heavy liquid, which does not. This term is smaller than in usual sampling calorimeter due to the smallness of the grains and the high density of the scintillator.
@@ 2 Experimental setup
The 16-channels GRAiNITA small-size prototype is shown in figure 1. It consists of a metallic container with an active volume of mm3 and it is equipped with 16 O-2(200) Kuraray WLS fibers (1 mm diameter) [8] arranged in square geometry and inter-spaced by 7 mm. In order to monitor the behaviour of the setup, a clear fiber, unpolished along 1 cm, is placed in the middle of the WLS ones and connected to a green LED222Broadcom HLMP-CM1A-560DD, wavelength from 500 to 540 nm. to inject the light in the active volume. Each of the 16 WLS fibers of GRAiNITA is coupled to a 1.3 1.3 mm2 SiPM S13360-1350E by Hamamastu [5] mounted on a PCB for supply, gain stability and amplification. The signals from the SiPMs are digitized at 3.2 GigaSample/s with a WaveCatcher module [3]. The number of photo-electrons is recorded in a 25 s windows corresponding to the scintillating decay time constant of the grains.
Data-taking took place during two days at the SPS beam (H2 line) delivering muons and pion secondary beams. Due to unstable beam conditions, the energy was varying but remained above few GeV, enough to ensure stable conditions for the measurements described in the following. In total about 2 million of muons events and 48 millions of pions events were recorded. Given the smallness of the GRAiNITA prototype, the estimated interaction probability for the pions is approximately 20%. The data acquisition was divided into two periods. In the first period (P1), ZnWO4 grains were immersed in water, while in the second period (P2) they were immersed in a water-based sodium polytungstate solution, hereafter referred to as the heavy liquid, characterized by a refractive index of 1.5 and a density of . For both periods, data were collected using pion and muon beams. Toward the end of the acquisition, partial leaks of the heavy liquid were observed due to a container closure not completely liquid-tight. Consequently, we chose to focus on the data obtained with water to evaluate non-uniformity effects, while using the heavy-liquid data primarily for calibration purposes, since no significant differences in non-uniformity are expected between the two liquid configurations. The experimental setup used is shown in figure 2 and consists of a set of 3 wire chambers (DWC) for reconstructing the beam tracks, 2 sets of dual scintillator for the triggering surrounding the GRAiNITA prototype.
@@ 3 Track reconstruction and checks
4 Assessments of the stochastic term of the energy resolution and of the light confinement
4.1 The stochastic term of the energy resolution
An encouraging number of approximately 400 photo-electrons was obtained with the cosmic muons [2] leading to a stochastic term on the energy resolution of the order of 1%. The much larger data sample recorded during the test beam and the possibility to better define the tracks used for the measurement calls for a new estimate. In order to avoid border effects, only the events where tracks are hitting the central region of the prototype are used, and the data recorded with muon beams is analysed. A Landau distribution convoluted with a Gaussian function as described in section 5.2 is fitted to the distribution of the total number of photo-electrons recorded on the 16 WLS fibres. The fit for the two periods is shown in figure 4.
The number of photoelectrons measured for water and for the heavy liquid are 479 and 452, respectively, and are in very good agreement with the values obtained using cosmic muons. [2].
@@ 4.2 Light confinement
Using the largest statistics run with pions from P2 and the four more central fibres, the average number of photo-electrons read by a given fibre as a function of the position of the track with respect to this fibre is shown in figure 5. These observations clearly confirm that the scintillation light remains confined near its point of production. The right-hand plot displays the dependence of the mean number of photo-electrons on the distance from the fibre centre. Approximately 50% of the detected light originates from particle interactions occurring within 2.7 mm from the fibre centre, corresponding to slightly less than half the spacing between adjacent fibres. These results and those described in section 3 are used to assess more precisely the fibre position. It is determined via the minimization of the sum of the number of photo-electrons in a square of bins in a bin windows around the expected position. The minimum in both directions is taken as the actual fibre position. The precision of this method is about the size of a bin (m).
@@ 6.3 Impact of the non-uniformity response on the constant term
A simplified model of a full scale calorimeter module is simulated using Geant4 Monte-Carlo [1]. The size of the calorimeter module is mm3. The depth of 400 mm corresponds to 25 radiation lengths (), which is sufficient to fully contain electromagnetic showers initiated by 25 GeV photons. This volume contains a grid of 24 by 24 WLS fibres evenly distributed in the corners of squares of 7 by 7 mm2, to reproduce the small-scale GRAiNITA prototype. The WLS fibres are simulated as cylinders of 1 mm diameter. In order to assess in a reasonable amount of CPU time the impact of the measured non-uniformity the grains are not simulated individually but a single material was simulated with partial contents of the scintillator and heavy liquid corresponding to the ones measured with the physical detector used for the beam test. The mass fractions of the detector material simulated are : 20.77 % of heavy liquid and 79.23 % of ZnWO4. The radiation length, calculated with the Geant4 framework, of this mixture of materials is 1.64 cm which aligns with manual cross-calculations.
@@ 7 Conclusion
Data from a two-day period in June 2024 using the H9 test beam of the SPS at CERN recorded with the small-scale GRAiNITA prototype is analyzed. These measurements confirm that the contribution to the energy resolution due to the number of photo-electrons is of the order of . With preliminary simulations of the grains and the heavy liquid [7], this term goes to about . More importantly, the data recorded during this test beam period enables a first estimation of the constant term associated with detector non-uniformity. Despite the limitations of these measurements due to the small size of the detector and the use of pions, the constant term due to the non-uniformity should be significantly below 1%.